a piece of iron2/3 yard long is cut into six equal pieces. how long is each piece in feet?
Answer by stanbon(75887)

a piece of iron2/3 yard long is cut into six equal pieces. how long is each piece in feet?
----
Note: (2/3) yard= (2/3)(3 ft) = 2 feet
Ans: 2/6 = 1/3 foot
Each piece is 1/3 foot long
=====================
Cheers,
Stan H.
=============